# Count-mean Sketch as an Optimized Framework for Frequency Estimation with Local Differential Privacy

### v1

@@ Appendix A Derivation of the Expectation
$E[\hat{y}^{(i)}[h_{j^{(i)}}(x)]]\\ =\sum_{j\in[k]}E\bigl{[}\mathbf{1}\{j=j^{(i)}\}R(h_{j}(X^{(i)}))[h_{j}(x)]\bigr{]}\\ =\sum_{j\in[k]}\frac{1}{k}c_{j}(X^{(i)},x),$ (25)
@@
Substituting $\forall x,x^{\prime}:\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}=\frac{1}{m}$ into the above equation, we have
@@
Appendix B Derivation of the Variance
@@
$Var\bigl{(}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{)}\\ =E\bigl{[}\bigl{(}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{)}^{2}\bigr{]}-\bar{c}(X^{(i)},x)^{2}$ (29)
where $E\bigl{[}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{]}=\bar{c}(X^{(i)},x)$ is proven in Eq. (25). Then,
@@
Therefore, when substituting Eq. (38) for $x^{\prime}\neq x$ and the aforementioned parameters with $x=x^{\prime}$ into Eq. (4), we have
$Var(\hat{f}(x))=\frac{m^{2}}{(m-1)^{2}n}\biggl{[}\\ \sum_{x^{\prime}\in[d]\backslash x}f(x^{\prime})\biggl{(}\bar{c}(x^{\prime},x)Var(R|=)+(1-\bar{c}(x^{\prime},x))\\ Var(R|\neq)+\bar{c}(x^{\prime},x)(1-\bar{c}(x^{\prime},x))\biggr{)}+f(x)Var(R|=)\biggr{]}.$ (39)
Thus, Eq. 5 is proved.
If $\bar{c}(x^{\prime},x)=\frac{1}{m}$, it becomes
$Var(\hat{f}(x))=\frac{m}{(m-1)^{2}n}\biggl{[}\\ \sum_{x^{\prime}\in[d]\backslash x}f(x)\biggl{(}Var(R|=)+(m-1)\\ Var(R|\neq)+\frac{1-m}{m}\biggr{)}+mf(x)Var(R|=)\biggr{]}\\ =\frac{m}{(m-1)^{2}n}\biggl{[}\bigl{(}1-f(x)\big{)}\biggl{(}Var(R|=)+(m-1)\\ Var(R|\neq)+\frac{1-m}{m}\biggr{)}+mf(x)Var(R|=)\biggr{]},$ (40)
which proves Eq. (6).
Appendix C Proof of Theorem 3.5
@@
Appendix D Proof regarding Optimized CMS
We first explore the optimized $m$ for the CMS+RR. Substituting Eq. (12) into Eq. (6), we have
Let’s first study the worst-case MSE of CMS+RR, which is equivalent to the worst-case variance $Var(\hat{f}(x)_{RR})$. Observe that the variance is linear to $f(x)$, we have $\max_{x}MSE(\hat{f}(x))=\max\{Var(\hat{f}(x)_{RR}|f(x)=0),Var(\hat{f}(x)_{RR}|f(x)=1)\}$. When $f(x)=1$, we have
$Var(\hat{f}(x)_{RR}|f(x)=0)=Var(\hat{f}(x)_{RR}|f(x)=1),$
we have $m=e^{\epsilon/2}+1$, and the corresponding variance is
$Var(\hat{f}(x)_{RR})=\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}.$
When $12\leq m<e^{\epsilon/2}+$ 1, we have $Var(\hat{f}(x)_{RR}|f(x)=1)<Var(\hat{f}(x)_{RR}|f(x)=0)$. Also, $Var(\hat{f}(x)_{RR}|f(x)=0)$ decreases with $m$ in this range. Therefore, when $12\leq m\leq e^{\epsilon/2}+$ 1, we have
That is, $m$ should never be smaller than $e^{\epsilon/2}+1$.
Define $f^{*}$ to be the upper bound of $f(x)$, i.e., $\forall x:f(x)\leq f^{*}$. When $e^{\epsilon/2}+1\leq m\leq e^{\epsilon}+1$, we have proven that $Var(\hat{f}(x)_{RR})$ increases with $f(x)$, so the goal becomes reducing $Var(\hat{f}(x)_{RR}|f(x)=f^{*})$. The optimized $m$ can be calculated by
$\frac{\partial Var(\hat{f}(x)_{RR}|f(x)=f^{*})}{\partial m}=0,$
which results in
$m=1+\frac{\Delta_{MSE}}{f^{*}e^{\epsilon}+1-f^{*}},$ (46)
where
The corresponding variance is
Notice that Eq. (46) is valid only when $e^{\epsilon/2}+1\leq m\leq e^{\epsilon}+1$, which translates to $f^{*}\leq\frac{1}{2}$. When $f^{*}>\frac{1}{2}$, we have $m<1+e^{\epsilon/2}$, which has been proven to be worse than $m=1+e^{\epsilon/2}$. Therefore, we have
which corresponds to
When it comes to symmetry and asymmetry RAPPOR, since their $Var(a)[a]$ and $Var(a)[b]$ are independent of $m$, the corresponding $Var(\hat{f}(x))$ decreases when $m$ increases. Thus, when $m$ is large enough, we have
@@ Appendix D Proof regarding Optimized CMS
For asymmetric RAPPOR, we realize $Var(\hat{f}(x)_{aRP})>Var^{*}(\hat{f}(x)_{RR})$ when $f^{*}>0$, so the CMS with asymmetric RAPPOR cannot serve as the optimized CMS. For symmetric RAPPOR, $Var(\hat{f}(x)_{sRP})>Var^{*}(\hat{f}(x)_{RR})$ when $0\leq f^{*}<\frac{1}{2}$, and $Var(\hat{f}(x)_{sRP})=Var^{*}(\hat{f}(x)_{RR})$ when $\frac{1}{2}\leq f^{*}\leq 1$. Thus, the CMS with symmetric RAPPOR is also ruled out. As a result, Theorem 15 is proved. Corollary 3.4 is proved automatically when substituting $f^{*}=1$ into Theorem 15.
We will first use RR to perturb the hashed values. Substituting Eq. (12) into Eq. (48), we obtain
$l_{2}(\hat{f}_{RR})=\frac{(d+e^{\epsilon}-1)m^{2}}{n(m-1)(e^{\epsilon}-1)^{2}}+\\ \frac{(d-1)(2m+e^{\epsilon}-1)}{n(m-1)(e^{\epsilon}-1)}.$ (49)
$m=1+\frac{\Delta_{l}}{e^{\epsilon}+d-1},$
where
In practice, we choose the closest integer to $1+\frac{\Delta_{l}}{e^{\epsilon}+d-1}$ as $m$. The corresponding minimum value is
$\min_{m}l_{2}(\hat{f}_{RR})=\frac{2(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}}.$ (50)
If the size of candidates $d$ is large enough, we have
$l_{2}(\hat{f}_{RR})\rightarrow\frac{4de^{\epsilon}}{n(e^{\epsilon}-1)^{2}},$ (51)
and the corresponding $m\rightarrow 1+e^{\epsilon}$.
Similar to the worst-case MSE, we could prove that the CMS with symmetric RAPPOR will always have larger $l_{2}$ loss than the CMS+RR. For asymmetric RAPPOR, its $l_{2}$ loss is the same as that of CMS+RR only when $d$ is large enough. However, its communication cost is too high because RAPPOR requires sending a vector of $m$ bits to the server, and Eq. (47) requires $m\rightarrow\infty$. To determine the necessary value of $m$, we solve the following equation for asymmetric RAPPOR:
@@
On the other hand, CMS+RR only requires $\log_{2}(1+e^{\epsilon})$ bits to output the perturbed result. Thus, CMS+RR is always preferred when optimizing $l_{2}$ loss due to its low communication cost. Now, we have proven the $l_{2}$-loss part in Theorem 3.7 and Corollary 3.5.
For $l_{1}$ loss, given the Cauchy-Schwarz inequality, we have $l_{1}\leq\sqrt{d\>l_{2}}$. Combining this with Eq. (50) leads to
$l_{1}(\hat{f})\leq\sqrt{\frac{2d(\Delta_{l}+de^{\epsilon})}{n(e^{\epsilon}-1)^{2}}}.$ (53)
Define $\lceil l_{1}\rceil$ to be the upper bound of $l_{1}$. If the size of candidates $d$ is large enough, we have
$\lceil l_{1}(\hat{f})\rceil\rightarrow\frac{2de^{\epsilon/2}}{\sqrt{n}e^{\epsilon}-1}.$ (54)
Similar to $l_{2}$ loss, the corresponding $m\rightarrow 1+e^{\epsilon}$ (in practice, we choose the closest integer). Thus, the $l_{1}$-loss part in Theorem 3.7 and Corollary 3.5 are proven.
Appendix E Proof regarding Imperfect Hashing
Let’s start by proving Theorem 3.8. Observe $\frac{\partial Var(g(x\mid\mathcal{H}^{\prime}))}{\partial m^{\prime}}$, and note that it is linear to $f(x)$. Considering the derivative at $f(x)=1$ and $f(x)=0$ yields:
$-\max\{\frac{2m^{\prime}Var(R|=)}{n(m^{\prime}-1)^{3}},\\ \frac{(m^{\prime}-1)(Var(R|\neq)+1)+(m^{\prime}+1)Var(R|=)}{n(m^{\prime}-1)^{3}}\}\\ \leq\frac{\partial Var(g(x|\mathcal{H}^{\prime}))}{\partial m^{\prime}}<0.$ (55)
@@ Appendix E Proof regarding Imperfect Hashing
$-\frac{2m^{\prime}Var(R|=)+m^{\prime}-1}{n(m^{\prime}-1)^{3}}\leq\frac{\partial Var(g(x|\mathcal{H}^{\prime}))}{\partial m^{\prime}}<0.$ (56)
At the same time, we have $m^{\prime}=m-\frac{m}{(2q+1)^{2}}$. Denote $Var(\hat{f}(x))$ as the variance of the CMS using perfect hashing. We want $Var(g(x|\mathcal{H}^{\prime}))$ to approach $Var(\hat{f}(x))$, which is formulated as $Var(g(x|\mathcal{H}^{\prime}))\leq Var(\hat{f}(x))(1+\tau)$, where $\tau$ is a small number like 0.01. This can be translated to
@@
thus proving Theorem 3.8. When $m=1+e^{\epsilon/2}$, $Var(\hat{f}(x))$ becomes $\frac{e^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$, and the left-hand side of Eq. (57) decreases with $\epsilon$. Thus, we can substitute $\epsilon=0$ into Eq. (57), and we have $2q+1>\sqrt{\frac{1}{\tau}}$. Corollary 3.6 is proved.
If $\forall x:Var(g(x))\leq(1+\tau)Var(\hat{f}(x))$, then $l_{2}(g)\leq(1+\tau)l_{2}(\hat{f}(x))$. When $m=1+e^{\epsilon}$, $Var(\hat{f}(x))$ increases with $f(x)$, so we have $Var(\hat{f}(x)|f(x)=0)\leq Var(\hat{f}(x))$. Thus, if
@@
is satisfied, Eq. (57) will also be satisfied. Substituting $\frac{4e^{\epsilon}}{(e^{\epsilon}-1)^{2}}$ into $Var(\hat{f}(x)|f(x)=0)$ in Eq. (58) and setting $m=1+e^{\epsilon}$, we also have the left-hand side of Eq. (58) decreasing with $\epsilon$. Thus, this inequality is still satisfied when $\epsilon=0$, which derives $2q+1>\sqrt{\frac{1}{\tau}}$. Thus, Corollary 3.7 is proved.
Now we will prove Theorem 3.9. First, by calculating the derivative of $Pr(h^{\prime}(x_{1})=h^{\prime}(x_{2}))$ with respect to $r$ to find the maximum, we have
$Pr(h^{\prime}(x_{1})=h^{\prime}(x_{2}))\leq\frac{1}{m}+\frac{1}{4mq(q+1)},$ (59)
which is equivalent to $\bar{c}(x,x^{\prime})\leq\frac{1}{m}+\frac{1}{4mq(q+1)}$.
Next, consider the $MSE(\hat{f}(x))$ in Eq. (7). We aim for $MSE(\hat{f}(x))$ to approach $Var(\hat{f}(x))$, where $\hat{f}(x)$ denotes the estimation with perfect pairwise independent hashing. Regarding the bias term of $MSE(\hat{f}(x))$, we have
$E[\hat{f}(x|\mathcal{H}^{\prime})]-f(x)=\frac{m}{m-1}(\bar{c}(x,x^{\prime})-\frac{1}{m})\sum_{x^{\prime}\in[d]\backslash x}f(x^{\prime})\\ \leq\frac{m}{m-1}\frac{1}{4mq(q+1)}=\frac{1}{4(m-1)q(q+1)}.$ (60)
Another term in $MSE(\hat{f}(x))$ is $Var(\hat{f}(x))$, which is linear to $\bar{c}(x,x^{\prime})$ and $\bar{c}(x,x^{\prime})^{2}$, and all the parameters of these two terms are positive. Therefore,
$\frac{Var[\hat{f}(x|\mathcal{H}^{\prime})]}{Var[\hat{f}(x)]}\leq\max\{\frac{\bar{c}(x,x^{\prime})}{1/m},\frac{\bar{c}(x,x^{\prime})^{2}}{1/m^{2}}\}\\ =\frac{(2q+1)^{4}}{16q^{2}(q+1)^{2}}.$ (61)
To satisfy
$MSE(f(x)|\mathcal{H}^{\prime})-MSE(\hat{f}(x))\leq 2\tau Var[\hat{f}(x)],$
it is equivalent to requiring
$(E[\hat{f}(x|\mathcal{H}^{\prime})]-f(x))^{2}<\tau Var[\hat{f}(x)]$
$\frac{Var[\hat{f}(x|\mathcal{H}^{\prime})]}{Var[\hat{f}(x)]}<1+\tau$
Substituting Eqs. (60) and (61) into the above equations proves Theorem 3.9.
Appendix F Proof regarding Correcting the Original CMS
F.1 Proof of Theorem 3.10
@@ Appendix F Proof regarding Correcting the Original CMS
F.2 Proof of Theorem 22
Given Markov’s inequality, Eq. (22) is equivalent to requiring
$\frac{\mathop{Var}_{\forall c_{j}(x,x^{\prime})}(E[\hat{f}(x)])}{\tau Var(\hat{f}(x))}\leq\delta.$ (64)
Theorem 3.10 has demonstrated that
$\mathop{Var}_{\forall c_{j}(x,x^{\prime})}(E[\hat{f}(x)])\leq\frac{1}{(m-1)k}$ (65)
The original CMS uses RAPPOR or Hadamard Encoding (HE) to perturb the hash values in Step 2, so their variance can be calculated using Eq. (5), which can be scaled down as
$Var(\hat{f}(x))>\frac{1}{n}\min\{Var(R|=),Var(R|\neq)\}\\ =\frac{Var(R|\neq)}{n},$ (66)
given that $Var(R|=)>Var(R|\neq)$ in both RAPPOR and HE. Combining Eq. (65) and (66), we have
$\frac{\mathop{Var}_{\forall c_{j}(x,x^{\prime})}(E[\hat{f}(x)])}{Var(\hat{f}(x))}<\frac{n}{(m-1)kVar(R|\neq)}.$
Substituting the above inequality into Eq. (64), the requirement of Eq. (22) is converted as
$\frac{n}{(m-1)kVar(R|\neq)}<\delta,$ (67)
which is the same as the condition of Theorem 22.
Appendix G Decoding Hadamard Encoding
@@
Appendix H Data Source of Section 5.1
We start with the accuracy. Since HE is equivalent to unbiased CMS+RR with $m=2$, we use Eq. (45) to compute its $\widehat{MSE}$, $l_{1}$ loss, and $l_{2}$ losses. The computed $\widehat{MSE}$ is the same as in [11]. The $l_{1}$ and $l_{2}$ losses of RHR were taken from [12], but it did not calculate the $\widehat{MSE}$. Section 4.3 shows that RHR is a skewed version of unbiased CMS, and the $\widehat{MSE}$ of RHR could be worse than that from the CMS+RR with $m=2$ (i.e., HE), so the $\widehat{MSE}$ of RHR is $\Omega\left(\left(\frac{e^{\epsilon}+1}{e^{\epsilon}-1}\right)^{2}\right)$. OLH is equivalent to the unbiased CMS with $m=1+e^{\epsilon}$, so we use Eq. (45) and follow Section 3.4 to calculate its $l_{1}$ loss, $l_{2}$ loss, and $\widehat{MSE}$. Section 4.5 shows that CMS+HE is equivalent to the unbiased CMS+RR with $m=2$ (i.e., HE), so its $l_{1}$ loss, $l_{2}$ loss, and $\widehat{MSE}$ are the same as HE. The $l_{1}$ and $l_{2}$ losses of SS are presented in [10]. Its $\widehat{MSE}$ is derived using Eq. (2) in [9] with
$p=\frac{ke^{\epsilon}}{ke^{\epsilon}+d-k}$
$k=d/(e^{\epsilon}+1)$
$f=1$
The $\widehat{MSE}$ of a-RP and s-RP is calculated using Eq. (2) in [9] as well. Both cases where $f(x)=0$ and $f(x)=1$ are calculated, and the worst-case variance is the maximum between them.
Next, we present the data source regarding communication cost. We first compute the cost of OCMS, which requires the server and a client to sync on the assigned hash function and the perturbed response $z^{(i)}$. Section 3.5 shows that a hash function is comprised of two random numbers, each ranging from 1 to $\Theta(d)$, so $2\log_{2}d$ bits are sufficient to represent one hash function. Then, we need another $\log_{2}m$ bits to sync the perturbed response $z^{(i)}$. Given that OCMS has either $m=1+e^{\epsilon/2}$ or $m=1+e^{\epsilon}$ depending on the optimization goal, the bits to transfer $z^{(i)}$ are close to $\epsilon/2$ and $\epsilon$, respectively. Thus, the overall communication cost of MSE-OCMS and $l$-OCMS is $22\log_{2}d+\epsilon/$ 2 and $2\log_{2}d+\epsilon$, respectively.
Appendix I Upper Bound of the worst-case MSE estimator
@@
Proof: Similar to Appendix C, we define $Y^{(i)}=\frac{m}{m-1}\hat{y}^{(i)}[h_{j}^{(i)}(x)]-\frac{1}{m-1}$, and $\frac{Y^{(i)}-A}{B-A}$ is a Bernoulli random variable (see Appendix C for the definition of $A$ and $B$). All the $X^{(i)}$ can be placed into two groups based on whether $X^{(i)}=x$. If $X^{(i)}=x$, $\frac{Y^{(i)}-A}{B-A}\sim\text{Ber}\left(\frac{e^{\epsilon}}{e^{\epsilon}+m-1}\right)$, where $\text{Ber}(p)$ denotes the Bernoulli distribution with probability $p$. On the other hand, if $X^{(i)}\neq x$, $\frac{Y^{(i)}-A}{B-A}\sim\text{Ber}\left(\frac{1}{m}\frac{e^{\epsilon}}{e^{\epsilon}+m-1}+\frac{m-1}{m}\frac{1}{e^{\epsilon}+m-1}\right)$. For convenience, we denote $p_{1}=\frac{e^{\epsilon}}{e^{\epsilon}+m-1}$ and $p_{2}=\frac{1}{m}\frac{e^{\epsilon}}{e^{\epsilon}+m-1}+\frac{m-1}{m}\frac{1}{e^{\epsilon}+m-1}$.

### v2

@@ Appendix D Derivation of the Expectation
$E[\hat{y}^{(i)}[h_{j^{(i)}}(x)]]=\sum_{j\in[k]}E\bigl{[}\mathbf{1}\{j=j^{(i)}\}\\ R(h_{j}(X^{(i)}))[h_{j}(x)]\bigr{]}=\sum_{j\in[k]}\frac{1}{k}c_{j}(X^{(i)},x),$ (24)
@@
Substituting $\forall x,x^{\prime}$, where $x\neq x^{\prime}$ and $\sum_{j\in[k]}\frac{c_{j}(x,x^{\prime})}{k}=\frac{1}{m}$ into the above equation, we have
@@
Appendix E Derivation of the Variance
@@
$Var\bigl{(}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{)}\\ =E\bigl{[}\bigl{(}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{)}^{2}\bigr{]}-\bar{c}(X^{(i)},x)^{2},$ (28)
where $E\bigl{[}\hat{y}^{(i)}[h_{j^{(i)}}(x)]\bigr{]}=\bar{c}(X^{(i)},x)$ given Eq. (24). Then,
@@
Therefore, when substituting Eq. (37) for $x^{\prime}\neq x$ and the aforementioned parameters with $x=x^{\prime}$ into Eq. (4), we derived Eq. (5). If $\bar{c}(x^{\prime},x)=\frac{1}{m}$, it becomes Eq. (6).
Appendix F Proof of Theorem 3.5
@@
Appendix G Proof regarding Preferring RR to RAPPOR in CMS
The variance of symmetry and asymmetry RAPPOR is listed below:
Since their $Var(a)[a]$ and $Var(a)[b]$ are independent of $m$, the corresponding $Var(\hat{f}(x))$ decreases when $m$ increases. Thus, when $m$ is large enough, we have
@@ Appendix G Proof regarding Preferring RR to RAPPOR in CMS
Following Appendix B and given unbiased CMS, the worst-case MSE of symmetry and asymmetry RAPPOR are identical to their variance, which are $\frac{e^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}$ and $\frac{f^{*}(e^{\epsilon}-1)^{2}+4^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$, respectively.
For asymmetric RAPPOR, we realize $Var(\hat{f}(x)_{aRP})>Var^{*}(\hat{f}(x)_{RR})$ when $f^{*}>0$, so the CMS with asymmetric RAPPOR cannot serve as the optimized CMS. For symmetric RAPPOR, $Var(\hat{f}(x)_{sRP})>Var^{*}(\hat{f}(x)_{RR})$ when $0\leq f^{*}<\frac{1}{2}$, and $Var(\hat{f}(x)_{sRP})=Var^{*}(\hat{f}(x)_{RR})$ when $\frac{1}{2}\leq f^{*}\leq 1$. Thus, the CMS with symmetric RAPPOR is also ruled out. As a result, The choice of RR is preferred.
Similar to Appendix B, we can use the variance to derive $l_{2}$ loss, which are $\frac{de^{\epsilon/2}}{n(e^{\epsilon/2}-1)^{2}}$ and $\frac{4de^{\epsilon}}{n(e^{\epsilon}-1)^{2}}$ for symmetric and asymmetric RAPPOR, respectively. we observe that CMS with symmetric RAPPOR will always have a larger $l_{2}$ loss than the CMS+RR. For asymmetric RAPPOR, its $l_{2}$ loss is the same as that of CMS+RR only when $d$ is large enough. However, its communication cost is too high because RAPPOR requires sending a vector of $m$ bits to the server, and Eq. (44) requires $m\rightarrow\infty$. To determine the necessary value of $m$, we solve the following equation for asymmetric RAPPOR:
@@
On the other hand, CMS+RR only requires $\log_{2}(1+e^{\epsilon})$ bits to output the perturbed result. Thus, CMS+RR is always preferred when optimizing $l_{2}$ loss due to its low communication cost.
Appendix H Proof regarding Imperfect Hashing
Let’s start by proving Theorem 3.9. Observe $\frac{\partial Var(g(x\mid\mathcal{H}_{api}))}{\partial m^{\prime}}$, and note that it is linear to $f(x)$. Considering the derivative at $f(x)=1$ and $f(x)=0$ yields:
$-\max\{\frac{2m^{\prime}Var(R|=)}{n(m^{\prime}-1)^{3}},\\ \frac{(m^{\prime}-1)(Var(R|\neq)+1)+(m^{\prime}+1)Var(R|=)}{n(m^{\prime}-1)^{3}}\}\\ \leq\frac{\partial Var(g(x|\mathcal{H}_{api}))}{\partial m^{\prime}}<0.$ (46)
@@ Appendix H Proof regarding Imperfect Hashing
$-\frac{2m^{\prime}Var(R|=)+m^{\prime}-1}{n(m^{\prime}-1)^{3}}\leq\frac{\partial Var(g(x|\mathcal{H}_{api}))}{\partial m^{\prime}}<0.$ (47)
At the same time, we have $m^{\prime}=m-\frac{m}{(2q+1)^{2}}$. Denote $Var(\hat{f}(x))$ as the variance of the CMS using perfect hashing. We want $Var(g(x|\mathcal{H}_{api}))$ to approach $Var(\hat{f}(x))$, which is formulated as $Var(g(x|\mathcal{H}_{api}))\leq Var(\hat{f}(x))(1+\tau)$, where $\tau$ is a small number like 0.01. This can be translated to
@@
thus proving Theorem 3.9. When $m=1+e^{\epsilon/2}$, $Var(\hat{f}(x))$ becomes $\frac{e^{\epsilon/2}}{(e^{\epsilon/2}-1)^{2}}$, and $\frac{Var(R|=)}{Var(\hat{f}(x))}$ decreases with $\epsilon$. Thus, we can substitute $\epsilon=0$ into Eq. (48), and we have $2q+1>\sqrt{\frac{1}{\tau}}$. The MSE part of Corollary 3.5 is proved.
If $\forall x:Var(g(x))\leq(1+\tau)Var(\hat{f}(x))$, then $l_{2}(g)\leq(1+\tau)l_{2}(\hat{f})$. When $m=1+e^{\epsilon}$, $Var(\hat{f}(x))$ increases with $f(x)$, so we have $Var(\hat{f}(x)|f(x)=0)\leq Var(\hat{f}(x))$. Thus, if
@@
is satisfied, Eq. (48) will also be satisfied. Substituting $\frac{4e^{\epsilon}}{(e^{\epsilon}-1)^{2}}$ into $Var(\hat{f}(x)|f(x)=0)$ in Eq. (49) and setting $m=1+e^{\epsilon}$, we also have the left-hand side of Eq. (49) decreasing with $\epsilon$. Since this inequality is still satisfied when $\epsilon=0$, it derives $2q+1>\sqrt{\frac{1}{\tau}}$. Thus, The $l_{2}$ part of Corollary 3.5 is proved.
Appendix I Proof of Theorem 3.10
@@
Appendix J Decoding Hadamard Encoding
@@
Appendix K Upper Bound of the worst-case MSE estimator
@@
Proof: Similar to Appendix F, we define $Y^{(i)}=\frac{m}{m-1}\hat{y}^{(i)}[h_{j}^{(i)}(x)]-\frac{1}{m-1}$, and $\frac{Y^{(i)}-A}{B-A}$ is a Bernoulli random variable (see Appendix F for the definition of $A$ and $B$). All the $X^{(i)}$ can be placed into two groups based on whether $X^{(i)}=x$. If $X^{(i)}=x$, $\frac{Y^{(i)}-A}{B-A}\sim\text{Ber}\left(\frac{e^{\epsilon}}{e^{\epsilon}+m-1}\right)$, where $\text{Ber}(p)$ denotes the Bernoulli distribution with probability $p$. On the other hand, if $X^{(i)}\neq x$, $\frac{Y^{(i)}-A}{B-A}\sim\text{Ber}\left(\frac{1}{m}\frac{e^{\epsilon}}{e^{\epsilon}+m-1}+\frac{m-1}{m}\frac{1}{e^{\epsilon}+m-1}\right)$. For convenience, we denote $p_{1}=\frac{e^{\epsilon}}{e^{\epsilon}+m-1}$ and $p_{2}=\frac{1}{m}\frac{e^{\epsilon}}{e^{\epsilon}+m-1}+\frac{m-1}{m}\frac{1}{e^{\epsilon}+m-1}$.